What simple interest rate will Susan need to secure to make $2,500 in interest on a $10,000 principal over 5 years?
Answer by checkley77(12844)

10000(1+x)^5=12500
10000(1+x)^5=12500
(1+x)^5=12500/10000
(1+x)^5=1.25
(1+x)=5thrt1.25
1+x=1.0456
x=1-1.0456
x=.0456 or 4.56% is the required interest rate.
Proof:
10000(1+.0456)^5=12500
10000(1.0456)^5=12500
10000*1.25=12500
12500=12500